### v1

# Active pattern formation emergent from single-species nonreciprocity

Abstract
Nonreciprocal forces violating Newton’s third law are common in a plethora of nonequilibrium situations ranging from predator-prey systems to the swarming of birds and effective colloidal interactions under flow. While many recent studies have focused on two species with nonreciprocal coupling, much less is examined for the basic single-component system breaking the actio and reactio equality of force within the same species. Here, we systematically derive the fundamental field theory of single-species nonreciprocal interactions from microscopic dynamics, leading to a generic framework termed Active Model N (N denoting nonreciprocity). We explore the rich dynamics of pattern formation in this intrinsic nonreciprocal system and the emergence of self-traveling states with persistent variation and flowing of active branched patterns. One particular new characteristic pattern is an interwoven self-knitting “yarn” structure with a unique feature of simultaneous development of micro- and bulk phase separations. The growth dynamics of a “ball-of-wool” active droplet towards these self-knitted yarn or branched states exhibits a crossover between different scaling behaviors. The mechanism underlying this distinct class of active phase separation is attributed to the interplay between force nonreciprocity and competition. Our predictions can be applied to various biological and artificial active matter systems controlled by single-species nonreciprocity.
Newton’s famous third law states that the force ${\mathbf{F}}_{1\rightarrow 2}$ which a body 1 exerts on another body 2 is, up to a sign, the same as the force $\mathbf{F}_{2\rightarrow 1}$ which body 2 exerts on body 1 such that there is equality of actio and reactio, i.e., $\mathbf{F}_{1\rightarrow 2}=-\mathbf{F}_{2\rightarrow 1}$. This law of reciprocity holds for all fundamental interactions and also for effective interactions in equilibrium situations. In nonequilibrium, however, reciprocity is typically broken such that $\mathbf{F}_{1\rightarrow 2}+\mathbf{F}_{2\rightarrow 1}\not=\mathbf{0}$ which implies that the center-of-mass of a two-body system is not force-free but experiences a driving force leading to spontaneous “active” motion of a pair.
There are two fundamentally different typical examples of nonreciprocity: either the two bodies are of different species (such as a predator trying to catch an escaping prey) or they are of the same type (such as two birds flying together where only the following bird can see the leading one). In the former case of two-species nonreciprocity typically only particles of different species couple nonreciprocally, while the latter case of single-species nonreciprocity concerns intrinsic nonreciprocal interactions within the same species. There are many examples for single-species nonreciprocity both for living systems such as birds, fish, bacteria [1], or starfish embryos [2], and artificial robotic and colloidal systems with vision-cone-like interactions [3]. Theory has considered various aspects of single-species nonreciprocity stemming from vision-cone effects [4, 5, 6, 7] or transverse interactions [1, 2, 8]. Systems governed by nonreciprocal interactions have shown potential technological importance, as demonstrated in various contexts ranging from robotic metamaterials [9, 10] to drug design [11].
In the field-theoretical description of active matter, most of the recent works on pattern formation induced by nonreciprocal interactions treat coupling-nonreciprocity of a binary system while ignoring nonreciprocity between the same species [12, 13, 14, 15, 16, 17]. These studies have provided crucial insights into the dynamics of multi-component systems [18, 19, 20] and a wealth of novel dynamical behavior has been discovered including binary active crystals [21, 22], symmetry-broken nonreciprocal phase transitions [16], as well as traveling bands [14] and transverse ripples [15].
Here we investigate the intrinsic case of active pattern formation induced by the original form of force nonreciprocity among single species. We start from the explicit form for the nonreciprocal interaction forces ${\mathbf{F}}_{2\rightarrow 1}$ and systematically derive a fundamental underlying continuum field theory based on microscopic dynamics, establishing a new model which we term Active Model N (where N denotes the novel aspect of nonreciprocity). A crucial advantage of this type of continuum field theories [23, 24, 25], in addition to the fact that they can reveal the complex mechanisms involved, is the accessibility to much larger spatial and temporal scales for a greater amount of particles as compared to discrete particle-based simulations. Unlike previous active field theories such as Active Model B [26] and its extensions [27, 28, 29], this model provides a generic framework to study the general class of single-species force nonreciprocity. It is a unifying field theory building an inherent connection between nonreciprocal interaction forces and self-propelled active matter, i.e., matter becoming active by nonreciprocity. We can then predict new types of self-traveling chiral patterns for which both spatial mirror symmetry and time-reversal symmetry are spontaneously broken. In particular, we discover a unique interwoven pattern reminiscent of self-knitting threads that we term “active yarn”, which self-organizes through simultaneous evolution of micro- and bulk phase separations over multiple length and time scales. This new class of active phase separation is induced by an intrinsic mechanism showing as the interplay between interparticle force nonreciprocity and the force competition that is beyond the traditional route, leading to the simultaneous occurrence of two distinct classes of collective behavior incorporating phase-separation-induced assembly and orientationally aligning, persistent collective self-migration. Notably, the size of the phase-separating yarn region is controllable (e.g., via the vision-cone angle). This mechanism has the potential for interesting applications in the realization of programmable matter, particularly in the form of controlled phase separation.
Fig. 1: Schematic of nonreciprocal single-species vision-cone model. An agent 1 with orientation $\mathbf{\hat{u}_{1}}$ sees and interacts with other agents within a vision cone of full opening angle $2\vartheta$. Agent 2 is included in the perception zone of agent 1, such that there is a resulting nonreciprocal force acting on agent 1 which is directed along $\mathbf{r}=\mathbf{r}_{2}-\mathbf{r}_{1}$. There is also a torque turning its orientation to be aligned with the orientation $\mathbf{\hat{u}_{2}}$ of the second agent. In this example, agent 2 with its own vision cone does not see agent 1 and thus there is no interaction between them that influences the motion of agent 2. Likewise, agent 3 is not visible for agent 1 such that there is no mutual interaction between agents 1 and 3 that influences the motion of either of them.
From microscopic to continuum field model
We start from a microscopic picture of particles or agents interacting through vision cones, as illustrated in Fig. 1. Consider a general form of pairwise nonreciprocal force between single-component particles with orientational degree of freedom on a plane, i.e.,
$\mathbf{F}_{2\rightarrow 1}=\hat{\mathbf{r}}f\left(|\mathbf{r}_{2}-\mathbf{r}_{1}|,\hat{\mathbf{u}}_{1}\cdot\hat{\mathbf{r}},\hat{\mathbf{u}}_{2}\cdot\hat{\mathbf{r}}\right),$ (1)
which satisfies both global translational symmetry via the dependence on $\mathbf{r}=\mathbf{r}_{2}-\mathbf{r}_{1}$ and global rotational symmetry via the dependence on $\hat{\mathbf{u}}_{i}\cdot\hat{\mathbf{r}}=\cos\varphi_{i}$, where $\hat{\mathbf{u}}_{i}$ denotes the unit vector of orientation of particle $i$ and $\hat{\mathbf{r}}=(\mathbf{r}_{2}-\mathbf{r}_{1})/|\mathbf{r}_{2}-\mathbf{r}_{1}|=(\mathbf{r}_{2}-\mathbf{r}_{1})/r$. The nonreciprocity of interaction is given by
$\mathbf{F}_{2\rightarrow 1}\neq-\mathbf{F}_{1\rightarrow 2}=-(-\hat{\mathbf{r}})f\left(r,-\hat{\mathbf{u}}_{2}\cdot\hat{\mathbf{r}},-\hat{\mathbf{u}}_{1}\cdot\hat{\mathbf{r}}\right),$ (2)
such that
$f(r,\cos\varphi_{1},\cos\varphi_{2})\neq f(r,-\cos\varphi_{2},-\cos\varphi_{1}).$ (3)
When particles interact within a finite vision cone of full opening angle $2\vartheta$ with $\vartheta\in(0,\pi]$ (Fig. 1), the above nonreciprocal condition can be achieved via
$f(r,\cos\varphi_{1},\cos\varphi_{2})=h(r)\Theta\left(-\cos\vartheta+\cos\varphi_{1}\right),$ (4)
giving $\mathbf{F}_{2\rightarrow 1}=\hat{\mathbf{r}}h(r)\Theta(-\cos\vartheta+\cos\varphi_{1})$ and $\mathbf{F}_{1\rightarrow 2}=-\hat{\mathbf{r}}h(r)\Theta(-\cos\vartheta-\cos\varphi_{2})$, with $h(r)>0$ for interparticle attraction and $h(r)<0$ for repulsion. The step function $\Theta$ is equal to one when the first particle can see the second one and zero otherwise, with nonreciprocity arising for $\vartheta\in(0,\pi)$.
Similarly, a nonreciprocal torque aligning the particle’s orientations can be set up as
$M(r,\hat{\mathbf{u}}_{1},\hat{\mathbf{u}}_{2})=\tau(r)\Delta\varphi_{21}\Theta\left(-\cos\vartheta+\cos\varphi_{1}\right),$ (5)
where $\Delta\varphi_{21}=\varphi_{2}-\varphi_{1}$ if $\varphi_{1}-\varphi_{2}\in[0,\pi]$ and $\Delta\varphi_{21}=\varphi_{2}-\varphi_{1}-2\pi$ if $\varphi_{1}-\varphi_{2}\in(\pi,2\pi)$, representing different directions of alignment rotation (a particle rotates in the direction closer to the desired orientation). The function forms of $h(r)$ and $\tau(r)$ depend on the specific types of attractive and repulsive forces and aligning torque.
To derive Active Model N, we start from the microscopic Langevin equations containing the nonreciprocal forces and torques. The dynamics of the full one-particle density field $\varrho(\mathbf{r},\hat{\mathbf{u}},t)$ is then described by the stochastically equivalent Smoluchowski equation. By a systematic expansion in terms of a particle density field $\rho(\mathbf{r},t)$ and a polarization density field $\mathbf{P}(\mathbf{r},t)$, the coarse-grained continuum field equations [24, 30] are obtained (see Methods and Supplementary Information). We then reach a new continuum density-field theory, i.e., Active Model N,
$\frac{\partial\rho}{\partial t}$ $=\nabla^{2}\rho+\bm{\nabla}\cdot\left[B_{1}\rho\mathbf{P}+B_{2}\rho\bm{\nabla}\rho+B_{3}\mathbf{P}\nabla^{2}\rho\right.$
$\quad\,+\left.2B_{3}\left(\mathbf{P}\cdot\bm{\nabla}\right)\bm{\nabla}\rho+B_{4}\rho\bm{\nabla}\nabla^{2}\rho\right],$ (6)
$\frac{\partial\mathbf{P}}{\partial t}$ $=\nabla^{2}\mathbf{P}-\tilde{D}_{\mathrm{R}}\mathbf{P}+2B_{1}\bm{\nabla}\rho^{2}+B_{1}\bm{\nabla}\cdot\left(\rho\underline{\mathbf{Q}}\right)$
$\quad\,-\tilde{D}_{\mathrm{R}}B_{5}\left(2\rho\mathbf{P}-\underline{\mathbf{Q}}\cdot\mathbf{P}\right),$ (7)
where the nematic order parameter tensor $\underline{\mathbf{Q}}$ is given by
$Q_{ij}$ $=B_{5}\Big{(}-P_{i}P_{j}+\frac{1}{2}\delta_{ij}P_{k}^{2}\Big{)}$
$\quad\,+\frac{B_{1}}{4\tilde{D}_{\mathrm{R}}}\left[\partial_{i}\left(\rho P_{j}\right)+\partial_{j}\left(\rho P_{i}\right)-\delta_{ij}\partial_{k}\left(\rho P_{k}\right)\right],$ (8)
with coupling coefficients $B_{i}$ ($i=1,...,5$) and the nondimensionalized rotational diffusion constant $\tilde{D}_{\mathrm{R}}$. Active Model N constitutes a minimal microscopic theory for single-species nonreciprocal interactions since its microscopic derivation (Supplementary Information) involves the minimal number of orders in the Fourier expansion [31] (namely one) required for obtaining a model that is not passive [24], and the minimal number of orders in the gradient expansion [32] (namely three) that are required for capturing the dynamics of active phase separation [23]. As shown in Methods, this model contains the previously developed Active Model B+ [28] as a limiting case, which is notable since Active Model B+ has been developed as a field theory for self-propelling particles [24].
The coupling coefficients $B_{i}$ can be explicitly calculated as moments of the microscopic nonreciprocal forces and torques. They vary with the vision-cone opening angle $\vartheta$ and the force and torque parameters. Detailed expressions are given in equations (18)–(22) of Methods for the choice
$h(r)=-F_{0}e^{-{r^{2}}/{a^{2}}}+F_{1}e^{-{r}/(\alpha a)},\quad\tau(r)=be^{-r/c},$ (9)
which describes a force with a short-ranged repulsive contribution of strength $F_{0}$ and a long-ranged attractive contribution of strength $F_{1}$, and a torque of strength $b$. The length scales $a$ and $c$ determine the range on which the forces and torques act. In the following we use this choice of soft interaction, and set weaker strength of attractive force (e.g., $F_{1}/F_{0}=10^{-3}$) and torque ($b/(F_{0}a)=0.3$) as compared to the repulsive force, but with longer interaction range ($\alpha=c/a=2$). Note that the nonreciprocal active field theory given by equations (6) and (7) is generic and the forms of model terms do not depend on the specific choices of force and torque functions. The combination of a short-ranged repulsive and a long-ranged attractive force holds for many interactions usually considered.
Bifurcations and phase diagram of active patterns
Two steady-state fixed points can be identified from the model equations (6) and (7), for a fully disordered state with constant particle density $\rho=\bar{\rho}$ and zero polarization density $\mathbf{P}=\bar{\mathbf{P}}=\mathbf{0}$, and a flocking state still with constant $\rho=\bar{\rho}$ but having nonzero fixed magnitude of polarization
$|\mathbf{P}|^{2}=|\bar{\mathbf{P}}|^{2}=-\frac{4}{B_{5}}\left(\bar{\rho}+\frac{1}{2B_{5}}\right),$ (10)
where $B_{5}$ is given by equation (22) (Methods), when
$1+2B_{5}\bar{\rho}<0,\quad{\rm i.e.,}\quad\bar{\rho}\vartheta>\frac{1}{4\pi\beta bc^{2}}.$ (11)
A primary bifurcation analysis can be conducted for the fully disordered state (Supplementary Information). For the soft interaction forms (9) chosen here, it yields a flocking transition (without phase separation) when equation (11) is satisfied at large enough average particle density $\bar{\rho}$ or opening angle $\vartheta$, with lower transition threshold for larger torque strength $b$ and range $c$, consistent with the torque aligning effect. The resulting homogeneous flocking phase is characterized by uniform distribution of particle density ($\rho=\bar{\rho}$) with constant polarization alignment ($\mathbf{P}=\bar{\mathbf{P}}\neq\mathbf{0}$), i.e., a uniform aligned phase where the orientational symmetry is spontaneously broken. More interesting transitions occur for the subsequent secondary bifurcation from this orientationally symmetry-breaking state, with the linearized dynamical equation for Fourier components of perturbations $\widehat{\rho}_{\mathbf{q}}$ and $\widehat{\mathbf{P}}_{\mathbf{q}}$ given by
$\frac{\partial}{\partial t}\left(\begin{array}[]{c}\widehat{\rho}_{\mathbf{q}}\\ \widehat{P}_{x\mathbf{q}}\\ \widehat{P}_{y\mathbf{q}}\end{array}\right)=\mathcal{L}(\mathbf{q},\bar{\rho},\bar{\mathbf{P}},\vartheta)\left(\begin{array}[]{c}\widehat{\rho}_{\mathbf{q}}\\ \widehat{P}_{x\mathbf{q}}\\ \widehat{P}_{y\mathbf{q}}\end{array}\right),$ (12)
where details of the dynamical matrix $\mathcal{L}$ are presented in Supplementary Information. When $\vartheta\neq\pi$ (with a limited vision perception $2\vartheta<2\pi$), $\mathcal{L}^{\dagger}\neq\mathcal{L}$, i.e., $\mathcal{L}$ is non-Hermitian, leading to nonreciprocal, parity-time (PT) symmetry breaking phase transitions [14, 16]. Note that here the system non-Hermiticity originates from vision-cone nonreciprocal interactions within single species, fundamentally different from the previous cases featured by asymmetric coupling between different populations (e.g., interspecies cross diffusivities) [15, 14, 17, 16] or different harmonic modes [16].
Fig. 2: Active patterns and phase diagram. a-e Simulation snapshots for various active patterns that emerge at average particle density $\bar{\rho}=1.5$, including active stripes at $\vartheta=30^{\circ}$ (a and Supplementary Video 1), active branches at $\vartheta=90^{\circ}$ (b and Supplementary Video 2), active yarn at $\vartheta=140^{\circ}$ (c and Supplementary Video 3) and $160^{\circ}$ (d and Supplementary Video 4), and a homogeneous flocking phase with vortex defects of polarization field $\mathbf{P}$ at the reciprocal limit $\vartheta=180^{\circ}$ (e and Supplementary Video 5). In a-d with patterns induced by nonreciprocity, the spatial profiles of particle density $\rho$ are presented, and the local distribution of vector field $\mathbf{P}$ are indicated as arrows in the insets which are enlarged portions of the boxed regions. The patterns self-travel opposite to the average direction of $\mathbf{P}$ (Supplementary Videos 1-5). In e the spatial profile of polarization magnitude $|\mathbf{P}|$ is shown. f The phase diagram of $\bar{\rho}$ vs $\vartheta$, with solid curves evaluated from the analytical results of bifurcation analysis and the symbols identified via outcomes of numerical simulations (those giving disordered or homogeneous phases are not shown).
We have calculated the corresponding $\bar{\rho}$ vs $\vartheta$ phase diagram, as shown in Fig. 2, by evaluating the analytical results of bifurcation analyses (Supplementary Information), with phase boundaries well agreeing with numerical simulations of the full model equations (6)–(8). A noteworthy phenomenon is the occurrence of phase transitions with the increase of $\vartheta$ or $\bar{\rho}$, i.e., from (i) the disordered state of both particle and polarization densities, to (ii) a traveling striped state of low- vs high-density microphase separation and local flocking alignment (Fig. 2a and Supplementary Video 1), then to (iii) homogeneous flocking with again disordered particle density, and back to (iv) a phase separated, flocking state with emergent dynamical chiral patterns (i.e., patterns for which spatial mirror symmetry is broken), including active branches (Fig. 2b and Supplementary Video 2) and active yarn pattern which shows as a phase consisting of traveling bands containing microphase-separated fine texture that are embedded in a homogeneous medium (see Fig. 2c,d and Supplementary Videos 3, 4), before (v) a reentrance to the homogeneous flocking phase at $\vartheta=\pi$. At the limit of $\vartheta=\pi$ with complete perception, the system becomes reciprocal and the dynamics of $\widehat{\rho}_{\mathbf{q}}$ and $\widehat{\mathbf{P}}_{\mathbf{q}}$ are decoupled, resulting in the stability of uniform flocking with homogeneous $\rho$ and the usual vector field dynamics for polarization $\mathbf{P}$ with motion and annihilation of vortex defects (Fig. 2e and Supplementary Video 5). The phase separation effects observed here are thus induced by vision-cone-type force nonreciprocity.
Although no explicit self-driving factor has been incorporated in the microscopic model, all phase-separated patterns self-travel unidirectionally, which leads to the breaking of time-reversal symmetry. The direction of pattern traveling is opposite to that of the average polarization field $\langle\mathbf{P}\rangle$, with average migration velocity $\bar{\mathbf{v}}\propto-\langle\mathbf{P}\rangle$. This is consistent with the microscopic picture: Due to nonreciprocity, each particle can only interact with the neighboring particles within its own vision cone centered around its orientation (Fig. 1). Since in the setup here the short-range repulsion is much stronger than the long-range attraction, the particle feels the net effect of repulsion from the particles in front of it (within the vision cone) but not from those behind, such that every particle ends up being pushed towards the direction opposite to its orientation. Given that $\mathbf{P}$ measures the particle orientation, collective particle migration occurs if equation (11) is satisfied with strong enough torque alignment. As particles are still being attracted to next neighbors, such a competing effect leads to the occurrence of phase separation at large enough particle density or opening angle. This phenomenon of nonreciprocity-induced phase separation with local orientational alignment then incorporates both types of collective behavior of active particles [33, 34], assembly via phase separation, and aligning collective self-migration, with mechanisms distinct from the previously known motility-induced [35] or nonreciprocal-torque-based [33] phase separation.
The unidirectional migration can be understood more quantitatively from the nonzero locally averaged force for a particle surrounded by its neighboring bath, i.e.,
[…]
Fig. 3: Dynamical processes of active pattern formation. Snapshots of density profile at different time stages of system evolution for two sample average densities $\bar{\rho}=0.7$ (a, c) and $\bar{\rho}=1$ (b, d), illustrating the rich dynamics of phase-separated fibers/strands which leads to the persistent variation of active branches (a, b with $\vartheta=110^{\circ}$, $90^{\circ}$; Supplementary Videos 6 and 7) or the self-knitting of active yarn (c, d with $\vartheta=130^{\circ}$; Supplementary Video 8).
Active branches and active yarn
One of our key findings is the emergence of self-migrating chiral patterns induced by single-species nonreciprocity, in the form of active branches and active yarn, at high enough opening angles or average particle densities. These patterns break time-reversal symmetry (traveling unidirectionally) and 2D parity symmetry (lacking mirror symmetry axes in space; Fig. 2b,c,d) for both particle and polarization density fields. The formation and nonlinear evolution of these exotic patterns are well beyond the stage of bifurcation and instability development, as illustrated in Fig. 3.
Figure 3 shows that two categories of pattern-forming dynamics can be identified. For intermediate vision-cone opening angles, we observe persistently varying active branches, while for large opening angles, the banding of active yarn occurs. For the dynamics of active branches there are two sub-types: At low enough average density (e.g., $\bar{\rho}=0.7$ for $\vartheta=110^{\circ}$; Supplementary Video 6), a microphase-separated stripe developed via instability sharpens to an individual fiber or thread, which splits and self-interweaves to form twisted or 2D double-helix type configurations. The strands then deform and bifurcate/branch, before breaking apart and constantly dispersing or regrouping. This double-twisting behavior does not occur at larger $\bar{\rho}$ (e.g., $\bar{\rho}=1$ for $\vartheta=90^{\circ}$; Supplementary Video 7), where one instead finds the curving and undulating of individual sharpened fibers during the secondary stage of phase separation. This is followed by their local self-braiding and bundling to form loose, varying clusters in the subsequent tertiary stage of evolution. For both sub-types the resulting branched traveling texture is highly dynamical, aggregating and dispersing incessantly.
For the second category – formation of active yarn at large opening angles – the branching occurs earlier and directly from the bent sharpening fibers. The multi-branched threads aggregate and merge to compact bands with fluctuating interfaces that separate from the homogeneous melt, yielding a self-knitting process of a strongly condensed, active yarn pattern (see Fig. 3 at $\vartheta=130^{\circ}$ and Supplementary Video 8). The pattern not only shows persistent variations of micro-textures within each band, but could also evolve via an alternating process of aggregation/breaking/rebanding with the change of traveling band orientation, while the overall pattern morphology still retains without coarsening with time (see Extended Data Fig. 1 and Supplementary Video 9).
Both active branches and active yarn patterns evolve at two different length and two separate time scales. The fine branched textures evolve rapidly and perpetually on short timescales, while the flow and variation of the entire pattern structure occur on longer timescales (see, e.g., Supplementary Videos 2-4, 9). As seen also in Extended Data Fig. 2 for the estimated probability distribution of velocity, the average migration velocity of the pattern, with value close to the sharp peak of the skewed distribution observed, corresponds to the overall timescale of pattern flow, while the extended range of the distribution, particularly for large magnitudes of velocity, reflects the fast variations of fine textures. Spatially, the local clustering during the variation of active branches, and more prominently the dense banding of active yarn, represent a large-scale phase separation between a surrounding homogeneous state and bundles or bands with fine-scale microphase-separated branched structures inside. Our simulations show that the larger length scale of bulk separation depends on the size of the simulation domain, suggesting that this is related to phase coexistence. The small length scale, on the other hand, depends on the physical parameters of the system and is about an order of magnitude smaller than the characteristic width of active yarn which is estimated in Extended Data Fig. 3. It is important to note that the formation of active yarn is beyond the traditional phase coexistence and is characteristic of a new dynamical pattern, since it involves the simultaneous development of small-scale multi-branching and larger-scale local bundling or banding (Fig. 3 and Supplementary Video 8), belonging to a different category of formation process as compared to active branches (noting also the difference of fine textures between active yarn with multi-branched strands within the highly condensed band and active branches consisting of mostly bifurcated fibers). Thus our results demonstrate the realizability of this unique type of active yarn pattern which self-organizes and evolves at multiple-scales through simultaneous microphase and bulk phase separations.
Microscopically, the banding phenomenon, particularly during the active yarn formation, can be understood by noting that, since the interaction combines a strong and short-ranged repulsive force with a weaker long-ranged attractive force (equation (9)), the interaction of a particle with its nearest neighbors is dominated by repulsion, whereas the interaction with its next-nearest neighbors is dominated by attraction. At very large vision-cone angle (e.g., $2\vartheta\geq 240^{\circ}$) with narrow blind spot, the net attraction from a broad range of next neighbors would be strong enough to forge and confine the particle, then bundling or banding together with other nearby particles while traveling concurrently. Thus narrower, denser bands can be expected for larger $\vartheta$ and smaller $\bar{\rho}$, consistent with the numerical simulation results shown in Extended Data Fig. 3. If the perception is further increased to $2\vartheta=2\pi$, there would be no blind spot and the interactions are reciprocal; the strong net repulsion would then push away particles from all directions ($\langle\mathbf{F}_{2\rightarrow 1}\rangle_{\rm local}=\mathbf{0}$ from equation (13)), making the density distribution disordered without any phase separation.
Active droplet growth dynamics
To further examine the chiral pattern dynamics, we simulate the growth process of an active droplet developing branched structures when embedded in a homogeneous flocking medium. As shown in Fig. 4, the self-traveling droplet domain grows slowly during the early-stage instability development of microphase separation to form weak stripes within, while the growth accelerates at intermediate time stage when the combination of microphase and bulk phase separations occurs during the secondary and tertiary processes of evolution; i.e., the stripes sharpen to become fibers which then curve/twist and branch, together with their local aggregation and bundling as well as segregating at a larger scale to the two internal ends of droplet, overall manifesting as a growing and moving “ball of wool”. A much faster growth takes place at later stage due to simulation boundary effects. These results are generic and consistent for different $\vartheta$ and $\bar{\rho}$ yielding active branched and yarn states.
Fig. 4: Active droplet growth. The droplet size $L(t)$ as a function of time $t$, at $\bar{\rho}=1$ and $\vartheta$ ranging from $90^{\circ}$ to $160^{\circ}$ for active branches or active yarn patterns (inset: $\bar{\rho}=1,1.2,1.5$ at $\vartheta=140^{\circ}$). The time-evolving snapshots, which are small portions of the large system simulated, correspond to the stars indicated on the $\vartheta=140^{\circ}$ curve. Also shown are the fittings to a power law scaling $L\propto t^{n}$ at early time (dashed lines), with $n=0.475\pm 0.001$ at $\vartheta=90^{\circ}$ and $n=0.435\pm 0.002$ at $\vartheta=140^{\circ}$, before the crossover to a linear regime of $L\propto t$.
It has been known [36] that the growth of droplet or domain size $L(t)$ obeys a power law behavior, $L\sim t^{n}$, with the scaling exponent $n=1/3$ for the diffusion-controlled mechanism (Lifshitz-Slyozov-Wagner law), $n=1/2$ for the curvature-driven growth, and $n=1$, i.e., linear growth, when the hydrodynamic advection effect dominates, while activity could possibly arrest phase separation and growth in “wet” systems [27, 37]. Our simulation results indicate a crossover between two scaling regimes before the influence of boundaries (Fig. 4), from an early-stage one showing only microphase separation with an exponent $n$ between $1/3$ and $1/2$, which can be attributed to a combination of diffusion-driven and curvature-driven effects, to a linear growth regime with $L\sim t$ when the growth mechanism is dominated by the combined phase separations induced by nonreciprocity. In contrast to the linear growth previously found in “wet” systems [27, 37], both advective effect and activity arrest are absent in the “dry” system studied here.
Conclusions
We have constructed a nonreciprocal continuum field theory, Active Model N, which incorporates the many-body effects induced by nonreciprocity and constitutes a minimal model for collective dynamics in the general class of single-species systems characterized by nonreciprocal forces. It allows us to reveal the intriguing results arising from the mechanism of competition between repulsive and attractive interactions of different strength and range and their combination with interparticle nonreciprocity and local alignment. This leads to the emergent phenomenon of single-species active pattern formation with PT symmetry breaking, a hallmark of nonreciprocal phase transitions. The resulting out-of-equilibrium chiral branched patterns, particularly the self-knitted active yarn, originate from a new class of active phase separation which simultaneously incorporates nonreciprocity-induced micro- and bulk phase separations, with local structural details being highly dynamical and varying persistently with time.
A similar behavior of banding and self-traveling of single-species agents with vision-cone interactions can be found in various real biological systems, such as the migration of caterpillars or ants [38] and the marching of penguins [39] or pedestrians [40]. The field theory developed here and our results of active pattern formation and dynamical mechanisms are expected to be generic and extendable to a range of nonreciprocal systems, particularly Janus colloids [3], that can be realized and controlled in experiments. Moreover, the active yarn phase, where a rather chaotic but patterned state is embedded within a homogeneous medium, has interesting potential applications in the realization of programmable materials. For example, if the particles are immersed in a fluid, this effect allows to achieve controlled mixing of this fluid in one region (the size of which can be controlled by tuning $\vartheta$ or $\bar{\rho}$) while leaving it unmixed elsewhere. If one additionally incorporates a mechanism in which the particles can adapt the value of $\vartheta$ in response to external stimuli, it would also allow for the realization of adaptive or even intelligent matter [41, 42].
Acknowledgements
We thank Filippo De Luca, Ken Elder, Helder Hugo, Jonas Mayer-Martins, and Simiso Mkhonta for helpful discussions. Z.-F.H. acknowledges support from the National Science Foundation under Grant No. DMR-2006446. M.t.V., R.W., and H.L. are funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-IDs 525063330, 433682494 – SFB 1459, and SPP 2265 – LO418/25.
Author contributions
Z.-F.H. and H.L. initiated and designed the research. M.t.V. conducted the theoretical derivation of continuum field equations. H.L., M.t.V., and R.W. developed the microscopic model. Z.-F.H. conducted the bifurcation analysis and performed numerical simulations. Z.-F.H., M.t.V., and H.L. wrote the manuscript. Z.-F.H., R.W., and H.L. prepared the figures. All authors analyzed the results, discussed the data, and revised the manuscript.
Competing interests
The authors declare no competing interests.
Extended Data Fig. 1: Time variation of active yarn pattern. During the nonreciprocity-induced self migration, the orientation and morphology of active yarn can vary with time, as illustrated in this sample simulation at $\bar{\rho}=1.5$ and $\vartheta=150^{\circ}$ (see also Supplemental Video 9).
Extended Data Fig. 2: Probability distribution of polarization and velocity. The probability density function (PDF) for the polarization component $P_{x}$ or $P_{y}$ is calculated from simulations at $\bar{\rho}=1.2$ and opening angle $\vartheta=90^{\circ}$ (active branches at $t=50000$) or $140^{\circ}$ (active yarn at $t=37500$, with $P_{x}=P_{\perp}$ and $P_{y}=P_{\parallel}$ in this example). It corresponds to the estimated probability distribution of locally coarse-grained particle velocity. The average migration velocity $\bar{\mathbf{v}}=(\bar{v}_{x},\bar{v}_{y})$, which is proportional to the reverse average polarization $-\langle\mathbf{P}\rangle$, is indicated in the inset images.
Extended Data Fig. 3: Width of active yarn as a function of opening angle $\vartheta$. Results are estimated from simulations at $\bar{\rho}=1.2$ and $1.5$ with system size $512\times 512$, showing the decrease of active yarn banding width with the increase of $\vartheta$ and the decrease of $\bar{\rho}$.